# Chapter 4 Discrete subgroups of semisimple Lie groups, beyond lattices

### 26 Feb 2024 v1

@@ 3.4 Higher Teichmüller theory
3.4.1 Towards a full list of higher Teichmüller spaces
@@
Without entering into technical details, let us mention briefly the role of Higgs bundles in this conjectural classification. See [33, 72] for details.
@@
Some of the connected components of the $G$-character variety of $\pi_{1}(S)$ can be distinguished using topological invariants. However, such invariants are not sufficient to distinguish them all in general. One fruitful approach is to use the fact, proved by Hitchin, that $(E,\varphi)\mapsto\|\varphi\|_{L^{2}(\Sigma)}^{2}$ defines a proper Morse function $f$ from $\mathcal{M}_{G}(\Sigma)$ to $\mathbb{R}_{\geq 0}$; therefore, the connected components of $\mathcal{M}_{G}(\Sigma)$ can be studied by examining the local minima of $f$. The zero locus $f^{-1}(0)$ of $f$ corresponds, in the $G$-character variety, to representations of $\pi_{1}(S)$ whose image lies in a compact subgroup of $G$; in particular, these representations are not injective and discrete, and so connected components for which $f$ has a local minimum of 00 cannot be higher Teichmüller spaces. This approach has already been successfully exploited to find and count almost all connected components of the $G$-character variety of $\pi_{1}(S)$ for simple $G$, including conjecturally all higher Teichmüller spaces: see [32, 33, 72].
3.4.2 Similarities with the classical Teichmüller space
@@
3.4.3 Higher higher Teichmüller spaces
@@
The second situation is in the context of pseudo-Riemannian hyperbolic geometry, which is by definition the study of pseudo-Riemannian manifolds (i.e. smooth manifolds with a smooth assignment, to each tangent space, of a nondegenerate quadratic form) which have constant negative sectional curvature. In signature $(p,q)$, such manifolds are locally modeled on the pseudo-Riemannian symmetric space $\mathbb{H}^{p,q}=\mathrm{PO}(p,q+1)/\mathrm{P}(\mathrm{O}(p)\times\mathrm{O}(q+1))$, which can be realised as an open set in projective space, namely $\{[v]\in\mathbb{P}(\mathbb{R}^{p+q+1})\,|\,\langle v,v\rangle_{p,q+1}<0\}$ where $\langle\cdot,\cdot\rangle_{p,q+1}$ is a symmetric bilinear form of signature $(p,q+1)$ on $\mathbb{R}^{p+q+1}$. For $q=0$ we recover the real hyperbolic space $\mathbb{H}^{p}$, with its projective model (1), and for $q=1$ the space $\mathbb{H}^{p,1}$ is the $(p+1)$-dimensional anti-de Sitter space (a Lorentzian analogue of the real hyperbolic space). Let $\Gamma_{0}=\pi_{1}(M)$ where $M$ is a closed hyperbolic $p$-manifold, with holonomy $\sigma_{0}:\Gamma_{0}\to\mathrm{O}(p,1)$, and let $\tau:\mathrm{O}(p,1)\hookrightarrow G=\mathrm{PO}(p,q+1)$ be the standard embedding. For $q=1$, Barbot [9] proved that the connected component of $\rho_{0}=\tau\circ\sigma_{0}$ in $\mathrm{Hom}(\Gamma_{0},G)$ consists entirely of injective and discrete representations (corresponding to holonomies of so-called globally hyperbolic spatially compact anti-de Sitter manifolds, studied in [112] for $p=2$). This was recently extended in [21] to general $p\geq 2$ and $q\geq 1$. In fact, the following more general result is proved in [21]: for $\Gamma_{0}=\pi_{1}(M)$ where $M$ is any closed topological manifold of dimension $p\geq 2$, the set of so-called $\mathbb{H}^{p,q}$-convex cocompact representations is a union of connected components in $\mathrm{Hom}(\Gamma_{0},G)$. These $\mathbb{H}^{p,q}$-convex cocompact representations are injective and discrete representations with a nice geometric behavior in $\mathbb{H}^{p,q}$ (see Section 5.2); they include the representations $\tau\circ\sigma_{0}:\Gamma_{0}=\pi_{1}(M)\to\mathrm{O}(p,1)\hookrightarrow G=\mathrm{PO}(p,q+1)$ above where $M$ is a closed hyperbolic manifold, but also other examples where $M$ can be quite “exotic” (see [108, 113] for $q=1$). These representations can have Zariski-dense image in $G$: see e.g. [21] for a bending argument as in Section 3.1.
@@
In these two situations, there are connected components in $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations, where $\Gamma_{0}$ is the fundamental group of an $n$-dimensional closed manifold with $n>2$ and $G$ is a semisimple Lie group with $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$. It is natural to call higher-dimensional higher-rank Teichmüller spaces (or higher higher Teichmüller spaces for short) the images of these components in the $G$-character variety of $\Gamma_{0}$. It would be interesting in the future to investigate whether these higher higher Teichmüller spaces have any topological or geometric analogies with classical Teichmüller space or its higher-rank counterparts, as above. See also [132, § 14] for some further discussion.
@@ 4 Classes of discrete subgroups in real rank one
4.1 Definitions
Consider, as in Section 2.1, the Riemannian symmetric space $X=G/K$, where $K$ is a maximal compact subgroup of $G$. If $G=\mathrm{SO}(n,1)$ (resp. $\mathrm{SU}(n,1)$, resp. $\mathrm{Sp}(n,1)$), then $X$ is the $n$-dimensional hyperbolic space over $\mathbb{R}$ (resp. $\mathbb{C}$, resp. the quaternions). If $G$ is the exceptional group $F_{4(-20)}$, then $X$ is the “hyperbolic plane over the octonions”.
There is a natural notion of convexity in $X$: any two points $x,y$ of $X$ are joined by a unique geodesic segment; we say that a subset $\mathcal{C}$ of $X$ is convex if this segment is contained in $\mathcal{C}$ for all $x,y\in X$. See [65, § 1.6] for more details. For any $\varepsilon>0$ and any subset $\mathcal{C}$ of $X$, we denote by $\mathcal{U}_{\varepsilon}(\mathcal{C})$ the uniform $\varepsilon$-neighbourhood of $\mathcal{C}$ in $X$.
@@ 4.1 Definitions
If there is a nonempty $\Gamma$-invariant convex subset $\mathcal{C}$ of $X$ such that $\Gamma\backslash\mathcal{C}$ is compact, then $\Gamma$ is automatically finitely generated, by the Švarc–Milnor lemma (see e.g. [64, Th. 8.37]). Thus the assumption that $\Gamma$ be finitely generated can be omitted in the definition of convex cocompactness. On the other hand, this assumption cannot be omitted in general in the definition of geometric finiteness: see [86].
@@
Bowditch [31] gave several equivalent definitions of geometric finiteness. Here we use a variation on his definition F5, where the uniform bound on the orders of finite subgroups of $\Gamma$ is replaced by the assumption that $\Gamma$ be finitely generated. The two definitions are equivalent by [31] and the Selberg lemma [123, Lem. 8].
@@
4.1.1 Limit sets and convex cores
Let $\partial_{\infty}X$ be the visual boundary of $X=G/K$, i.e. the set of equivalence classes of geodesic rays in $X$ for the equivalence relation “to remain at bounded distance”. There is a natural topology on $\overline{X}:=X\sqcup\partial_{\infty}X$ that extends that of $X$ and makes $\overline{X}$ compact, and the action of $G$ on $X$ extends continuously to $\overline{X}$ (see e.g. [65, § 1.7]). For instance, as in Sections 3.1 and 3.4, if $G=\mathrm{PO}(n,1)$ and $X=\mathbb{H}^{n}$, then we can realise $X$ as the open subset (1) of $\mathbb{P}(\mathbb{R}^{n+1})$ where some quadratic form of signature $(n,1)$ is negative, and $\overline{X}$ is then the closed subset of $\mathbb{P}(\mathbb{R}^{n+1})$ where the quadratic form is nonpositive, endowed with the topology from $\mathbb{P}(\mathbb{R}^{n+1})$ and the natural action of $G=\mathrm{PO}(n,1)$.
@@ 4.2.2 Deformations
On the other hand, geometric finiteness is in general not stable under small deformations. If one restricts to small deformations that are cusp-preserving (i.e. that keep parabolic elements parabolic), then stability holds for $G=\mathrm{PO}(n,1)$ when $n\leq 3$ or when all cusps have rank $\geq n-2$, but not in general. See e.g. [80, App. B] for more details and references.
4.2.3 Homomorphisms
@@ 4.3 Examples
    On the other hand, for $G=\mathrm{PO}(n,1)$ with $n\geq 3$, there exist finitely generated discrete subgroups of $G$ which are not geometrically finite. The first examples were constructed by Bers for $n=3$ (“singly degenerate” Kleinian groups, for which the domain of discontinuity $\Omega_{\Gamma}$ is simply connected): see [93, § 2].
  • Any discrete subgroup of $G=\mathrm{PO}(n,1)$ generated by the orthogonal reflections in the faces of a finite-sided right-angled polyhedron of $\mathbb{H}^{n}$ is geometrically finite; it is convex cocompact if and only if no distinct facets of the polyhedron have closures meeting in $\partial_{\infty}\mathbb{H}^{n}$ (see [59, § 4]).
  • The Schottky groups of Section 3.1 are geometrically finite; the strong Schottky groups (for which $B_{1}^{\pm},\dots,B_{m}^{\pm}$ have pairwise disjoint closures) are convex cocompact. Their limit sets are Cantor sets. The set $\Omega$ of Remark 4.7 is the domain of discontinuity $\Omega_{\Gamma}=\partial_{\infty}X\smallsetminus\Lambda_{\Gamma}$ of $\Gamma$ in $\partial_{\infty}X$ from Section 4.2.
  • Any quasi-Fuchsian group $\Gamma=\rho(\pi_{1}(S))$ as in Section 3.1 is convex cocompact. The limit set $\Lambda_{\Gamma}$ is a topological circle in $\partial_{\infty}\mathbb{H}^{3}$ (see Figure 3). The quotient $\Gamma\backslash\mathcal{C}^{\mathsf{cor}}_{\Gamma}$ is homeomorphic to $S\times[0,1]$.
  • The small deformations of cocompact lattices of $G^{\prime}=\mathrm{SO}(n,1)$ inside $G=\mathrm{SO}(n+1,1)$ from Section 3.1 are convex cocompact by Facts 4.18 and 4.19 (see also Remark 4.8.(2)).
4.4 A few characterisations of convex cocompactness
4.4.1 Preliminaries
@@
As in Section 2.3, a group $\Gamma_{0}$ is called Gromov hyperbolic if it is finitely generated and acts properly discontinuously, by isometries, with compact quotient, on some Gromov hyperbolic proper geodesic metric space $Y$; in that case, we can take $Y$ to be $\mathrm{Cay}(\Gamma_{0})$. As in the proof of Theorem 4.11, the Gromov boundary $\partial_{\infty}\Gamma_{0}$ of $\Gamma_{0}$ is then the visual boundary of $Y$, endowed with the action of $\Gamma_{0}$ extending that on $Y$. The Gromov boundary $\partial_{\infty}\Gamma_{0}$ does not depend on $Y$ up to $\Gamma_{0}$-equivariant homeomorphism. An important property is that the action of $\Gamma_{0}$ on $\partial_{\infty}\Gamma_{0}$ is a convergence action: for any sequence $(\gamma_{k})_{k\in\mathbb{N}}$ of pairwise distinct elements of $\Gamma_{0}$, up to passing to a subsequence, there exist $w^{+},w^{-}\in\partial_{\infty}\Gamma_{0}$ such that $\gamma_{k}\cdot w\to w^{+}$ for all $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$, uniformly on compact sets. Moreover, any infinite-order element of $\Gamma_{0}$ has two fixed points in $\partial_{\infty}\Gamma_{0}$, one attracting and one repelling. The group $\Gamma_{0}$ is called elementary if it is finite (in which case $\partial_{\infty}\Gamma_{0}$ is empty) or if it admits a finite-index subgroup which is cyclic (in which case $\partial_{\infty}\Gamma_{0}$ consists of two points). If $\Gamma_{0}$ is not elementary, then the set of attracting fixed points of infinite-order elements of $\Gamma_{0}$ is infinite and dense in $\partial_{\infty}\Gamma_{0}$. See e.g. [14] for details.
@@
Finally, we denote by $\mathtt{d}_{X}$ the metric on the Riemannian symmetric space $X=G/K$ (see Section 2.1). We fix a basepoint $x_{0}\in X$, and a Riemannian metric $\mathtt{d}_{\partial_{\infty}X}$ on the visual boundary $\partial_{\infty}X$.
4.4.2 A few classical characterisations
Many interesting characterisations of convex cocompactness have been found by various authors, including Beardon, Bowditch, Maskit, Sullivan, Thurston, Tukia, and others. We now give a few. We refer to [93, 94] for more details and references, as well as further characterisations (e.g. in terms of conical limit points). We also refer to [31, 93] for characterisations of geometric finiteness.
@@ 4.4.3 Sketches of proofs
We may assume that the basepoint $x_{0}$ belongs to the convex core $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$. By the Švarc–Milnor lemma (see e.g. [64, Th. 8.37]), if $\Gamma_{0}$ acts properly discontinuously, by isometries, with compact quotient, on a proper geodesic metric space $M$, then $\Gamma_{0}$ is finitely generated and any orbital map $\gamma\mapsto\gamma\cdot m$ is a quasi-isometric embedding: there exist $c,c^{\prime}>0$ such that $\mathtt{d}_{M}(m,\gamma\cdot m)\geq c\,\mathtt{d}_{\mathrm{Cay}(\Gamma_{0})}(e,\gamma)-c^{\prime}$ for all $\gamma\in\Gamma_{0}$. We apply this to the convex core $M=\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, endowed with the restriction of the metric $\mathtt{d}_{X}$. ∎
@@
The orbital map $\gamma\mapsto\rho(\gamma)\cdot x_{0}$ from $\Gamma_{0}$ to $X$ extends to a map from $\mathrm{Cay}(\Gamma_{0})$ to $X$ sending edges of $\mathrm{Cay}(\Gamma_{0})$ to geodesic segments of $X$. The fact that $\rho$ is a quasi-isometric embedding implies the existence of $c,c^{\prime}>\nolinebreak 0$ such that any geodesic of $\mathrm{Cay}(\Gamma_{0})$ is sent to a $(c,c^{\prime})$-quasigeodesic in $X$, and the Morse lemma (see e.g. [64, Th. 11.40 & 11.105]) states that $(c,c^{\prime})$-quasigeodesics are uniformly close to actual geodesics in $X$. Therefore the orbit $\rho(\Gamma_{0})\cdot x_{0}$ is quasiconvex: there exists a uniform neighbourhood $\mathcal{U}$ of $\rho(\Gamma_{0})\cdot x_{0}$ in $X$ such that any geodesic segment between two points of $\rho(\Gamma_{0})\cdot x_{0}$ is contained in $\mathcal{U}$. We conclude using the fact (see [31, Prop. 2.5.4]) that any quasiconvex subset of $X$ lies at finite Hausdorff distance from its convex hull in $X$. ∎
@@
    for $M=\mathrm{Cay}(\Gamma_{0})$, the inequality (6) is “almost” an equality: $\mathrm{length}^{\infty}_{\mathrm{Cay}(\Gamma_{0})}(g)\geq\mathrm{transl}_{\mathrm{Cay}(\Gamma_{0})}(g)-8\delta$ where $\delta\geq 0$ (hyperbolicity constant) is such that all triangles of $\mathrm{Cay}(\Gamma_{0})$ are $\delta$-thin.
(Then (4) holds with $c^{\prime\prime}=8\delta$.) We note that actually
@@
Indeed, (ii) is based on the fact that $X$ is a CAT(0) space: any geodesic triangle of $X$ is “at least as thin” as a triangle with the same side lengths in the Euclidean plane. Applying this to a geodesic triangle with vertices $m,g\cdot m,g^{2}\cdot m$, we see that if $m^{\prime}$ is the midpoint of the geodesic segment $[m,g\cdot m]$ (so that $g\cdot m^{\prime}$ is the midpoint of $[g\cdot m,g^{2}\cdot m]$), then $\mathtt{d}_{X}(m^{\prime},g\cdot m^{\prime})\leq\mathtt{d}_{X}(m,g^{2}\cdot m)/2$. By induction on $k$, we obtain that for any $m\in M$ and any $k\geq 1$, there exists $m_{k}\in M$ such that $\mathtt{d}_{X}(m,g^{2^{k}}\cdot m)\geq 2^{k}\,\mathtt{d}_{X}(m_{k},g\cdot m_{k})\geq 2^{k}\,\mathrm{transl}_{M}(g)$. We conclude by dividing by $2^{k}$ and passing to the limit.
(i) can be proved in a similar way, replacing the CAT(0) inequality $\mathtt{d}_{X}(m^{\prime},g\cdot m^{\prime})\leq\mathtt{d}_{X}(m,g^{2}\cdot m)/2$ by the Gromov hyperbolicity inequality $\mathtt{d}_{X}(m^{\prime},g\cdot m^{\prime})\leq\mathtt{d}_{X}(m,g^{2}\cdot m)/2+4\delta$ (see [50, Prop. 5.1]). ∎

### 28 Mar 2024 v2

@@ 4 Higher Teichmüller theory
Towards a full list of higher Teichmüller spaces
@@
Without entering into technical details, let us mention briefly the role of Higgs bundles in this conjectural classification. See [32, 34, 74] for details.
@@
Some of the connected components of the $G$-character variety of $\pi_{1}(S)$ can be distinguished using topological invariants. However, such invariants are not sufficient to distinguish them all in general. One fruitful approach is to use the fact, proved by Hitchin, that $(E,\varphi)\mapsto\|\varphi\|_{L^{2}(\Sigma)}^{2}$ defines a proper Morse function $f$ from $\mathcal{M}_{G}(\Sigma)$ to $\mathbb{R}_{\geq 0}$; therefore, the connected components of $\mathcal{M}_{G}(\Sigma)$ can be studied by examining the local minima of $f$. The zero locus $f^{-1}(0)$ of $f$ corresponds, in the $G$-character variety, to representations of $\pi_{1}(S)$ whose image lies in a compact subgroup of $G$; in particular, these representations are not injective and discrete, and so connected components for which $f$ has a local minimum of 00 cannot be higher Teichmüller spaces. This approach has already been successfully exploited to find and count almost all connected components of the $G$-character variety of $\pi_{1}(S)$ for simple $G$, including conjecturally all higher Teichmüller spaces: see [32, 33, 34, 74].
Similarities with the classical Teichmüller space
@@
Higher higher Teichmüller spaces
@@
The second situation is in the context of pseudo-Riemannian hyperbolic geometry, which is by definition the study of pseudo-Riemannian manifolds (i.e. smooth manifolds with a smooth assignment, to each tangent space, of a nondegenerate quadratic form) which have constant negative sectional curvature. In signature $(p,q)$, such manifolds are locally modeled on the pseudo-Riemannian symmetric space $\mathbb{H}^{p,q}=\mathrm{PO}(p,q+1)/\mathrm{P}(\mathrm{O}(p)\times\mathrm{O}(q+1))$, which can be realised as an open set in projective space, namely $\{[v]\in\mathbb{P}(\mathbb{R}^{p+q+1})\,|\,\langle v,v\rangle_{p,q+1}<0\}$ where $\langle\cdot,\cdot\rangle_{p,q+1}$ is a symmetric bilinear form of signature $(p,q+1)$ on $\mathbb{R}^{p+q+1}$. For $q=0$ we recover the real hyperbolic space $\mathbb{H}^{p}$, with its projective model (1), and for $q=1$ the space $\mathbb{H}^{p,1}$ is the $(p+1)$-dimensional anti-de Sitter space (a Lorentzian analogue of the real hyperbolic space). Let $\Gamma_{0}=\pi_{1}(M)$ where $M$ is a closed hyperbolic $p$-manifold, with holonomy $\sigma_{0}:\Gamma_{0}\to\mathrm{O}(p,1)$, and let $\tau:\mathrm{O}(p,1)\hookrightarrow G=\mathrm{PO}(p,q+1)$ be the standard embedding. For $q=1$, Barbot [9] proved that the connected component of $\rho_{0}=\tau\circ\sigma_{0}$ in $\mathrm{Hom}(\Gamma_{0},G)$ consists entirely of injective and discrete representations (corresponding to holonomies of so-called globally hyperbolic spatially compact anti-de Sitter manifolds, studied in [114] for $p=2$). This was recently extended in [21] to general $p\geq 2$ and $q\geq 1$. In fact, the following more general result is proved in [21]: for $\Gamma_{0}=\pi_{1}(M)$ where $M$ is any closed topological manifold of dimension $p\geq 2$, the set of so-called $\mathbb{H}^{p,q}$-convex cocompact representations is a union of connected components in $\mathrm{Hom}(\Gamma_{0},G)$. These $\mathbb{H}^{p,q}$-convex cocompact representations are injective and discrete representations with a nice geometric behaviour in $\mathbb{H}^{p,q}$ (see Section 2); they include the representations $\tau\circ\sigma_{0}:\Gamma_{0}=\pi_{1}(M)\to\mathrm{O}(p,1)\hookrightarrow G=\mathrm{PO}(p,q+1)$ above where $M$ is a closed hyperbolic manifold, but also other examples where $M$ can be quite “exotic” (see [110, 115] for $q=1$). These representations can have Zariski-dense image in $G$: see e.g. [21] for a bending argument as in Section 1.
@@
In these two situations, there are connected components in $\mathrm{Hom}(\Gamma_{0},G)$ consisting entirely of injective and discrete representations, where $\Gamma_{0}$ is the fundamental group of an $n$-dimensional closed manifold with $n>2$ and $G$ is a semisimple Lie group with $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$. It is natural to call higher-dimensional higher-rank Teichmüller spaces (or higher higher Teichmüller spaces for short) the images of these components in the $G$-character variety of $\Gamma_{0}$. It would be interesting in the future to investigate whether these higher higher Teichmüller spaces have any topological or geometric analogies with classical Teichmüller space or its higher-rank counterparts, as above. See also [134, § 14] for some further discussion.
@@ 4 Classes of discrete subgroups in real rank one
1 Definitions
Consider, as in Section 1, the Riemannian symmetric space $X=G/K$, where $K$ is a maximal compact subgroup of $G$. If $G=\mathrm{SO}(n,1)$ (resp. $\mathrm{SU}(n,1)$, resp. $\mathrm{Sp}(n,1)$), then $X$ is the $n$-dimensional hyperbolic space over $\mathbb{R}$ (resp. $\mathbb{C}$, resp. the quaternions). If $G$ is the exceptional group $F_{4(-20)}$, then $X$ is the “hyperbolic plane over the octonions”.
There is a natural notion of convexity in $X$: any two points $x,y$ of $X$ are joined by a unique geodesic segment; we say that a subset $\mathcal{C}$ of $X$ is convex if this segment is contained in $\mathcal{C}$ for all $x,y\in X$. See [66, § 1.6] for more details. For any $\varepsilon>0$ and any subset $\mathcal{C}$ of $X$, we denote by $\mathcal{U}_{\varepsilon}(\mathcal{C})$ the uniform $\varepsilon$-neighbourhood of $\mathcal{C}$ in $X$.
@@ 1 Definitions
If there is a nonempty $\Gamma$-invariant convex subset $\mathcal{C}$ of $X$ such that $\Gamma\backslash\mathcal{C}$ is compact, then $\Gamma$ is automatically finitely generated, by the Švarc–Milnor lemma (see e.g. [65, Th. 8.37]). Thus the assumption that $\Gamma$ be finitely generated can be omitted in the definition of convex cocompactness. On the other hand, this assumption cannot be omitted in general in the definition of geometric finiteness: see [88].
@@
Bowditch [31] gave several equivalent definitions of geometric finiteness. Here we use a variation on his definition F5, where the uniform bound on the orders of finite subgroups of $\Gamma$ is replaced by the assumption that $\Gamma$ be finitely generated. The two definitions are equivalent by [31] and the Selberg lemma [125, Lem. 8].
@@
Limit sets and convex cores
Let $\partial_{\infty}X$ be the visual boundary of $X=G/K$, i.e. the set of equivalence classes of geodesic rays in $X$ for the equivalence relation “to remain at bounded distance”. There is a natural topology on $\overline{X}:=X\sqcup\partial_{\infty}X$ that extends that of $X$ and makes $\overline{X}$ compact, and the action of $G$ on $X$ extends continuously to $\overline{X}$ (see e.g. [66, § 1.7]). For instance, as in Sections 1 and 4, if $G=\mathrm{PO}(n,1)$ and $X=\mathbb{H}^{n}$, then we can realise $X$ as the open subset (1) of $\mathbb{P}(\mathbb{R}^{n+1})$ where some quadratic form of signature $(n,1)$ is negative, and $\overline{X}$ is then the closed subset of $\mathbb{P}(\mathbb{R}^{n+1})$ where the quadratic form is nonpositive, endowed with the topology from $\mathbb{P}(\mathbb{R}^{n+1})$ and the natural action of $G=\mathrm{PO}(n,1)$.
@@ 2 Properties
On the other hand, geometric finiteness is in general not stable under small deformations. If one restricts to small deformations that are cusp-preserving (i.e. that keep parabolic elements parabolic), then stability holds for $G=\mathrm{PO}(n,1)$ when $n\leq 3$ or when all cusps have rank $\geq n-2$, but not in general. See e.g. [82, App. B] for more details and references.
Homomorphisms
@@ 3 Examples
    On the other hand, for $G=\mathrm{PO}(n,1)$ with $n\geq\nolinebreak 3$, there exist finitely generated discrete subgroups of $G$ which are not geometrically finite. The first examples were given by Bers for $n=3$ (“singly degenerate” Kleinian groups, for which the domain of discontinuity $\Omega_{\Gamma}$ is simply connected): see [95, § 2].
  • Any discrete subgroup of $G=\mathrm{PO}(n,1)$ generated by the orthogonal reflections in the faces of a finite-sided right-angled polyhedron of $\mathbb{H}^{n}$ is geometrically finite; it is convex cocompact if and only if no distinct facets of the polyhedron have closures meeting in $\partial_{\infty}\mathbb{H}^{n}$ (see [60, § 4]).
  • The Schottky groups of Section 1 are geometrically finite; the strong Schottky groups (for which $B_{1}^{\pm},\dots,B_{m}^{\pm}$ have pairwise disjoint closures) are convex cocompact. Their limit sets are Cantor sets. The set $\Omega$ of Remark 3.3 is the domain of discontinuity $\Omega_{\Gamma}=\partial_{\infty}X\smallsetminus\Lambda_{\Gamma}$ of $\Gamma$ in $\partial_{\infty}X$ from Section 2.
  • Any quasi-Fuchsian group $\Gamma=\rho(\pi_{1}(S))$ as in Section 1 is convex cocompact. The limit set $\Lambda_{\Gamma}$ is a topological circle in $\partial_{\infty}\mathbb{H}^{3}$ (see Figure 3). The quotient $\Gamma\backslash\mathcal{C}^{\mathsf{cor}}_{\Gamma}$ is homeomorphic to $S\times[0,1]$.
  • The small deformations of cocompact lattices of $G^{\prime}=\mathrm{SO}(n,1)$ inside $G=\mathrm{SO}(n+1,1)$ from Section 1 are convex cocompact by Facts 4.7 and 4.8 (see also Remark 3.4.(2)).
4 A few characterisations of convex cocompactness
Preliminaries
@@ 4 A few characterisations of convex cocompactness
As in Section 3, a group $\Gamma_{0}$ is called Gromov hyperbolic if it is finitely generated and acts properly discontinuously, by isometries, with compact quotient, on some Gromov hyperbolic proper geodesic metric space $Y$; in that case, we can take $Y$ to be $\mathrm{Cay}(\Gamma_{0})$. As in the proof of Theorem 3.7, the Gromov boundary $\partial_{\infty}\Gamma_{0}$ of $\Gamma_{0}$ is then the visual boundary of $Y$, endowed with the action of $\Gamma_{0}$ extending that on $Y$. The Gromov boundary $\partial_{\infty}\Gamma_{0}$ does not depend on $Y$ up to $\Gamma_{0}$-equivariant homeomorphism. An important property is that the action of $\Gamma_{0}$ on $\partial_{\infty}\Gamma_{0}$ is a convergence action: for any sequence $(\gamma_{k})_{k\in\mathbb{N}}$ of pairwise distinct elements of $\Gamma_{0}$, up to passing to a subsequence, there exist $w^{+},w^{-}\in\partial_{\infty}\Gamma_{0}$ such that $\gamma_{k}\cdot w\to w^{+}$ for all $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$, uniformly on compact sets. Moreover, any infinite-order element of $\Gamma_{0}$ has two fixed points in $\partial_{\infty}\Gamma_{0}$, one attracting and one repelling. The group $\Gamma_{0}$ is called elementary if it is finite (in which case $\partial_{\infty}\Gamma_{0}$ is empty) or if it admits a finite-index subgroup which is cyclic (in which case $\partial_{\infty}\Gamma_{0}$ consists of two points). If $\Gamma_{0}$ is not elementary, then the set of attracting fixed points of infinite-order elements of $\Gamma_{0}$ is infinite and dense in $\partial_{\infty}\Gamma_{0}$. See e.g. [14] for details.
@@
Finally, we denote by $\mathtt{d}_{X}$ the metric on the Riemannian symmetric space $X=G/K$ (see Section 1). We fix a basepoint $x_{0}\in X$, and a Riemannian metric $\mathtt{d}_{\partial_{\infty}X}$ on the visual boundary $\partial_{\infty}X$.
A few classical characterisations
Many interesting characterisations of convex cocompactness have been found by various authors including Beardon, Bowditch, Maskit, Sullivan, Thurston, Tukia, and others. We now give a few. We refer to [95, 96] for more details and references, as well as further characterisations (e.g. in terms of conical limit points). We also refer to [31, 95] for characterisations of geometric finiteness.
@@
We may assume that the basepoint $x_{0}$ belongs to the convex core $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$. By the Švarc–Milnor lemma (see e.g. [65, Th. 8.37]), if $\Gamma_{0}$ acts properly discontinuously, by isometries, with compact quotient, on a proper geodesic metric space $M$, then $\Gamma_{0}$ is finitely generated and any orbital map $\gamma\mapsto\gamma\cdot m$ is a quasi-isometric embedding: there exist $c,c^{\prime}>0$ such that $\mathtt{d}_{M}(m,\gamma\cdot m)\geq c\,\mathtt{d}_{\mathrm{Cay}(\Gamma_{0})}(e,\gamma)-c^{\prime}$ for all $\gamma\in\Gamma_{0}$. We apply this to the convex core $M=\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, endowed with the restriction of the metric $\mathtt{d}_{X}$. ∎
@@
The orbital map $\gamma\mapsto\rho(\gamma)\cdot x_{0}$ from $\Gamma_{0}$ to $X$ extends to a map from $\mathrm{Cay}(\Gamma_{0})$ to $X$ sending edges of $\mathrm{Cay}(\Gamma_{0})$ to geodesic segments of $X$. The fact that $\rho$ is a quasi-isometric embedding implies the existence of $c,c^{\prime}>\nolinebreak 0$ such that any geodesic of $\mathrm{Cay}(\Gamma_{0})$ is sent to a $(c,c^{\prime})$-quasigeodesic in $X$, and the Morse lemma (see e.g. [65, Th. 11.40 & 11.105]) states that $(c,c^{\prime})$-quasigeodesics are uniformly close to actual geodesics in $X$. Therefore the orbit $\rho(\Gamma_{0})\cdot x_{0}$ is quasiconvex: there exists a uniform neighbourhood $\mathcal{U}$ of $\rho(\Gamma_{0})\cdot x_{0}$ in $X$ such that any geodesic segment between two points of $\rho(\Gamma_{0})\cdot x_{0}$ is contained in $\mathcal{U}$. We conclude using the fact (see [31, Prop. 2.5.4]) that any quasiconvex subset of $X$ lies at finite Hausdorff distance from its convex hull in $X$. ∎
@@
    for $M=\mathrm{Cay}(\Gamma_{0})$, the inequality (6) is “almost” an equality: $\mathrm{length}^{\infty}_{\mathrm{Cay}(\Gamma_{0})}(g)\geq\mathrm{transl}_{\mathrm{Cay}(\Gamma_{0})}(g)-8\delta$ where $\delta\geq 0$ is a hyperbolicity constant for $\mathrm{Cay}(\Gamma_{0})$ (i.e. all triangles of $\mathrm{Cay}(\Gamma_{0})$ are $\delta$-thin).
Indeed, then (4) will hold with $c^{\prime\prime}=8\delta$. We note that actually
@@
[…]
(i) can be proved in a similar way, replacing the CAT(0) inequality $\mathtt{d}_{X}(m^{\prime},g\cdot m^{\prime})\leq\mathtt{d}_{X}(m,g^{2}\cdot m)/2$ by the Gromov hyperbolicity inequality $\mathtt{d}_{X}(m^{\prime},g\cdot m^{\prime})\leq\mathtt{d}_{X}(m,g^{2}\cdot m)/2+4\delta$ (see [51, Ch. 10, Prop. 5.1]). ∎